What is the name of the polygon whose interior angles have a sum of 540 degrees?

Thanks,
Amanda
: What is the name of the polygon whose interior angles have a sum of 540 degrees?

Thanks,
Amanda

Answer by Fombitz(1756)

The formula for the sum of interior angles for an n-sided polygon is
S=180*(n-2)
In your case,
540=180*(n-2)
n-2=3
n=5
A five sided polygon is a pentagon.
Question 149616: What is the name of the polygon whose interior angles have a sum of 540 degrees?

Thanks,
Amanda
: What is the name of the polygon whose interior angles have a sum of 540 degrees?

Thanks,
Amanda

Answer by ayman11(2)

( n - 2 ) * 180 = 540

n - 2 = 540 ÷ 180

n - 2 = 3
n = 2 + 3
n = 5
then its called a " Penagon "